FEL Upgrade Project Weekly Brief
January 6-10, 2003

Happy New Year to all of our readers.  The workforce returned partially on Jan. 2 and in full force on Jan. 6th
after a much needed break in the hectic installation and commissioning activities.  This week we made good
progress in the areas of photogun commissioning, SRF commissioning of the injector cryounit, installation of 2 of
the 5 magnets in the first recirculation arc, qualification of the last two families of dipole magnets, and qualification of
the high reflector optical cavity.

Commissioning Progress:
Details are given in the WBS sections given below.  A summary is as follows:
1) A high temperature bakeout of the photogun/light box assembly was completed over the holidays.  This week
we conditioned the internal NEG pumps and cesitator.  On Monday we will begin an in situ cathode heat clean,
attempt to make a photoactive cathode surface and measure the photo response.
2) Final SRF commissioning of the injector cryounit started on Wednesday evening and continues tonight.  One of
the two cavities (#3) is fully commissioning and we are working on the second cavity.
3) The 180 dipole magnet and vacuum chamber for the first arc was installed; the first reverse bend dipole (GQ)
for this arc was also installed this week
4) The difficult and lengthy process of qualifying the first of the GQ and GX family of magnets was completed this
week. Qualification of subsequent members of each family is now a modest, (several day) procedure. A work
round was needed in the initial GX qualification as noted the WBS9 section.
5) The vacuum and mirror motion assembly of the high reflector optical cavity was qualified this week

Today (Jan. 10th), the Under Secretary of Energy, Robert Card, visited Jefferson Lab for the first time.  The visit
included a briefing on all the lab’s programs including a complete facility tour. Prior to visiting the lab today, Mr.
Card visited Northrop Grumman-Newport News. During this visit the NG-NN management kindly mentioned the
long term collaboration between JLab and NG-NN on FEL development and applications.

The weekly financial report was prepared and forwarded to the DOE and DOD program offices.  Given the
extremely tight FY03 budget situation, all budgetary transactions are being closely monitored.

A trip has been set up next week for visit on Jan. 16th to the Air Force Research Laboratory.
John Eric, our AF Program Manager will host a visit and seminar by F.Dylla and H. Helvajian on the FEL Upgrade
project and Aerospace Corporation’s development of a laser microengineering station for the FEL Facility

WBS 4 (Injector):
The photogun, light box and downstream beamline was baked out to 250C until Dec. 24th.
During the cooldown we noticed that another one of the downstream beam position monitor feedthroughs had
opened up a small leak during the bakeout or cooldowon.  The leak was repaired temporarily until the next vacuum
cycle so that we could go on with gun commissioning.  The bakeout had done a reasonable job of converting the
residual hydrorocarbon contamination we observed prior to the bakeout to methane. The achieved total pressure
levels after the bakeout was in the 1 x10-10 range with hydrogen dominating and methane in the 3-4 x 10 –11
range. This about an order of magnitude higher than we like to be for cathode production and lifetime but we
should be able to make a starting cathode. Proceeding in that direction we conditioned the in-situ NEG pumps in
the gun and light box chambers and slowly degassed the cesiators.  Early next week we will begin the process of
preparing a photoactive cathode by heat cleaning followed by NF3/Cs deposition.  If we get a sufficient
photoresponse, then we will apply high voltage and slowly raise the voltage to operating levels
(clamped at 350k V for now).

Gun HVPS - Ready to be closed for operations.  It has been left open for I&C testing of their interface box.  The
Conditioning/Running Resistor is ready for installation.

WBS 6 (RF):
Quarter HVPS - SRF testing of cavity 3 was conducted this week.  The HVPS for cavity 4 developed a problem,
but has been repaired.  Cavity 4 is scheduled for testing this evening.

Zone 2 - Completed and ready for operations.

Zone 4 - Ready for additional SRF testing this evening.

WBS 8 (Instrumentation):
The gun high voltage power supply digital control card arrived, was assembled and tested. Thanks to EECAD for
their support.  The EECAD package for the digital I/O module has been signed off and returned to document
control.  The analog control card will be complete next week, there is a hand wired version that we will begin the
run with.  EECAD has completed the following VME front panel fabrication drawings:
F0150, F0163, and F0174. Check plots have been received for the 30Hz variable timing box, the LSS User
Shutter Control Box Fabrication, the LSS Lab Interface Box Fabrication and the IPPS Main Control Board

Magnet Thermal Interlock (klixon) cables in Zones 5 and 1 have been terminated and the low level checkout
completed successfully.  Rack terminations for the 2nd Arc, Wiggler, dispersion and injection shunts continue.
Installation of the 2 BeamViewer interface boxes for the ultra viewers was completed allowing local control.
ITV0F02 (Ceramic Viewer) was installed after the gun bake and high level checkout performed successfully.  Low
level checkout of the Wiggler beamviewers (ITV4F06A-D) was also completed successfully.

The hardware and software for operating the optical cavities is nearly complete.  The problems with LVDT position
sensors was the gain in the VME board, this was calibrated and all works now.  Five stepper motor front panel
modules were populated in house.  The first complete 8 channel stepper motor chassis was completed, tested,
checked out and installed.  This will support the Optics Group to complete the optical cavity testing.  The balance
of the stepper drive chassis only lack the front panels.  Stepper motor system wiring and HVPS Interlock system
wiring diagrams have been turned over to EECAD for check prints.  Progress has also been made in completing
the drive laser LVDT's and Pico motor terminations in the Injector.

About 1/2 of the carpet padding in the penetrations in the vault has been replaced with fire retardant foam blocks.
More blocks are on order to complete the replacement.  We still need to verify the tightness following the
replacement.  The current TSOP for ODH in the vault expires on January 25th.  An extension is in routing to
extend this until mid-April.  We will need to perform a complete ODH assessment prior to the expiration of this

WBS 9 (Transport):
Work Arounds
Substitution of IR Demo 180° Bend (DY) for the Upgrade 180° Bend (GY) in the First Arc.
• The DY magnet was mounted on its stands in the enclosure.
 Diversion of the beam after the Wiggler.
• The vacuum group is working on the vacuum pipes in the extraction area beyond the wiggler.
Optical Chicane Dipoles (GW)
• Magnet Test will do testing on these UV Line magnets after the IR Magnets are fully tested.
Injector Dipoles (DU/DV)
• No action
Arc 180 Degree Dipoles (GY)
• Wang NMR completed the first GY Magnet and shipped it on Monday.  It will arrive next Monday.
Arc Bend & Reverse Bend Dipoles (GQ, GX)
• The first GQ Dipole was mounted in the first arc.
• The data on the first GX for IR operation is in evaluation.  We moved the centerline toward the narrow side by
0.73 inches, using the wedge shape to reducer the excessive magnet length.  This also eliminated any angular error.
However this movement is enough to hit the edge of the start of the field region.  David Douglas is having Chris
Tennant, his graduate student, look at the data along arcs at the inner edge and outer edge of the required good
field region in order to more closely mimic actual beam path.

 The first GQ and this GX match core fields within spec at all momentums and integrals at 80 and 145 MeV/c.  An
additional shim added to the pole face shim should match the integral at 210.

• In shoving over the magnet as above, there is only one interference with the beam chamber.  The optical cavity
pipe interferes with the small triangular post of steel left on the other side of the hole in the back leg.  In response,
we had the machine shop remove the post all together because it would be too narrow for maintain structural
integrity after being trimmed down.
Trim Quad (QT)
• The trim coil to create the GC vertical corrector function within the trim quad was awarded to Milhous Control
and is in fabrication.
Sextupole (SF)
• No progress
Octupole (OT)
• The procurement process is stopped at the requisition stage because of lack of funds
Beam Line and Vacuum
• Only one chamber remains to be fabricated (by our shop) for the entire machine that is the last chamber in Optical
Chicane.  The chambers for the Bend-Reverse Bend line in both Arcs as well as the first three in the Optical chine
are all qualified by the vacuum group.  Next week is planned to insert the chambers in the magnets.  The two 180*
Bend chambers await the receipt of the diagnostic windows from Ceramaseal to receive qualification by the
vacuum group.
• The vacuum Group has started up again at installing the various spool pieces to complete the vacuum through the

WBS 11 (Optics):
During the holiday shutdown we continued to pump down the high reflector assembly, and monitored its progress
remotely.  The pressure vs time data, coupled with analyses of the RGA spectra, showed that we had a large
amount of water adsorbed on the metal and Kapton-insulated wire.  Increasing the pumping speed, removing all
unnecessary wiring, and a 40 hour hot dry nitrogen gas purge reduced the total pressure by an order of magnitude
and reduced the percentage of water by two orders of magnitude.  This was sufficient to operate the leadscrew
under vacuum and thus answered two questions: (1) does the Vespel leadscrew nut self-lubricate the leadscrew
while under vacuum, and does it perform well, and (2) test the cooling loop conveyor chain.    Tests this week
answered both questions.  After very slowly driving the mirror assembly to the top of it's travel, the friction had
reduced to the point that subsequent upward travel could be accomplished at ~ 20x higher speeds.  In fact, our
upward speed was limited by the motor and possibly motion control parameters.  Downward travel improved as
well.  At this time, we can do a complete traverse in the downward direction, corresponding to changing the laser
state from low power, broadband operation to short wavelength, high power operation, in less than 30 minutes.
We can traverse in the opposite direction in less than 10 minutes.  At no time during our cycling did the cooling
loop conveyor chain begin to bind.  So this component was successfully tested as well.  We then vented the
chamber with dry nitrogen and checked for particulates.  We did not find any visible particulates in the bottom of
the vacuum vessel.  However, we found some on the top of the "C" assembly and foundation plate.  This appears
to be due to the dry lubricant on the rotary feedthrough bearing spalling off.  We can minimize the dispersal by
designing a "cup" to capture the material.  We also plan to work with the vendor to determine what they used and
see if we can get a more robust coating.  We now turn to completing the internal wiring and water cooling loop for
the high reflector, and then do the same wiring for the outcoupler.  Fabrication of the diagnostic viewer/laser safety
shutter and the insertable mirrors continues and we anticipate receipt of parts the third week of January.

Other activities:
The drive laser developed another small leak towards the end of December.  We determined the cause and fixed
it.  We will finish realigning the laser through the optics in the clean room late today or early Monday.  We are
delayed because the air conditioner for the clean room failed yesterday.  Hardware for the JTO-funded drive laser
project continues to arrive.  We assisted in delivering two laser systems (one is IR tunable, ps pulsewidth, the other
fixed wavelength fs pulsewidth) acquired by Profs. Gunter  Luepke and Michael Kelley, respectively.  They will be
installed in User Lab 6.